# RMP-YOLO: A Robust Motion Predictor for Partially Observable Scenarios even if You Only Look Once

### v1

@@ Abstract
We introduce RMP-YOLO, a unified framework designed to provide robust motion predictions even with incomplete input data. Our key insight stems from the observation that complete and reliable historical trajectory data plays a pivotal role in ensuring accurate motion prediction. Therefore, we propose a new paradigm that prioritizes the reconstruction of intact historical trajectories before feeding them into the prediction modules. Our approach introduces a novel scene tokenization module to enhance the extraction and fusion of spatial and temporal features. Following this, our proposed recovery module reconstructs agents’ incomplete historical trajectories by leveraging local map topology and interactions with nearby agents. The reconstructed, clean historical data is then integrated into the downstream prediction modules. Our framework is able to effectively handle missing data of varying lengths and remains robust against observation noise, while maintaining high prediction accuracy. Furthermore, our recovery module is compatible with existing prediction models, ensuring seamless integration. Extensive experiments validate the effectiveness of our approach, and deployment in real-world autonomous vehicles confirms its practical utility. In the 2024 Waymo Motion Prediction Competition, our method, RMP-YOLO, achieves state-of-the-art performance, securing third place.
@@ I INTRODUCTION
    Our method won third place in the 2024 Waymo Motion Prediction Competition. We deploy our algorithms on real vehicles to validate the effectiveness of our methods. We will open-source our code to support the robotics community after the paper is accepted.
@@ II-A Motion Prediction
Recent advancements[7, 8, 9, 10, 11] in multi-agent motion prediction have introduced several innovative approaches aimed at improving prediction accuracy and efficiency. Among these, two of the most influential prediction paradigms are MTR [7] and QCNet [8]. MTR addresses the trade-off between goal-based prediction[12, 13, 14, 15] and direct regression-based[16, 11, 14, 17, 18] prediction, by its proposed motion query pairs, while QCNet proposes a symmetric[19], translation-invarnt[20] input representation which enables reusable stream prediction. Plenty derivative works have expanded on these foundational approaches, including EDA [21], MTR++ [9], ControlMTR [22], SIMPL [16], MGTR [23], and others [24, 25, 26, 27, 28]. However, a common limitation shared by these methods is that they are typically trained on nearly complete input trajectories. As a result, their performance may degrade when confronted with imperfect or incomplete data during inference (see Fig.7). To address this, we propose the RMP-YOLO framework, building upon MTR, to effectively handle input data imperfections while maintaining high prediction accuracy.
@@ II-B Imperfect Data Recovery
For feature distillation method[32], a well-trained teacher model using complete input trajectories guides a student model using incomplete input trajectories. POP[29] combines both methods to address partially observable prediction. While effective, this approach is complex, computationally expensive, and relies heavily on the quality of the teacher model. Some studies [33, 34] have investigated enhancing the robustness of predictors against input noise from the perspective of data poisoning. However, these approaches often overlook the impact of incomplete trajectories. Unlike previous methods, we propose a unified framework that explicitly recovers clean historical trajectories in the encoder stage and integrates this new information back into the agent tokens for further information fusion. As a result, our approach is robust to both noisy and incomplete data while maintaining excellent prediction performance. Additionally, our proposed recovery module is lightweight and plug-and-play.
@@ III-A Problem Formulation and Input Representation
Following MTR, we adopt an vectorized representation for both maps and agents. And our focus is on marginal motion prediction. The historical trajectories of ${N_{a}}$ traffic participants are denoted as $\mathcal{A}=\{a_{1},a_{2},\dots,a_{N_{a}}\}$, where each agent $a_{i}\in\mathbb{R}^{T_{p}\times F_{1}}$ has $T_{p}$ past timesteps and $F_{1}$ feature dimensions. The corresponding map is partitioned into $N_{l}$ polylines =$\mathcal{M}=\{m_{1},m_{2},\dots,m_{N_{l}}\}$, with each polyline $m_{i}\in\mathbb{R}^{N_{p}\times F_{2}}$ comprising $N_{p}$ points and $F_{2}$ feature dimensions. The predictor will anticipate $K$ different modality future trajectories $\mathcal{Z}=\{z_{1},z_{2},\dots,z_{N_{a}}\}$ over the future $T_{f}$ timesteps, where $z_{i}=\{z_{i}^{1},z_{i}^{2},\dots,z_{i}^{K}\}\in\mathbb{R}^{K\times T_{f}\times D}$. The confidence score for $z_{i}$ are denoted as $p_{i}=\{p_{i}^{1},p_{i}^{2},\dots,p_{i}^{K}\}$. Then for selected agent $z_{i}$, existing motion prediction task aims to estimate the distribution:
@@ III-B Network Encoder
where $\text{MHA}(\cdot,\cdot,\cdot)$ represents the multi-head attention function, $Q^{0}=[\mathcal{M}_{map},\mathcal{A}_{agent}]\in\mathbb{R}^{(N_{a}+N_{m})\times D_{1}}$, and $\mathcal{K}(\cdot)$ denotes the $K$-nearest neighbours (KNN) algorithm, which is used to identify the $K$ nearest tokens relative to each query. The term $\text{PE}(\cdot)$ refers to the sinusoidal positional encoding assigned to input tokens, incorporating the most recent position of each agent and the central point of each map polyline. The final output of the local attention layer will be snet into the MTR decoder, $[\mathcal{M}_{map},\mathcal{A}_{agent}]=Q^{Final}$.
@@ IV Experiments
TABLE I: Prediction on the test leaderboard of the motion prediction track of the Waymo Open Dataset Challenge. Our approach is termed as RMP, i.e., Robust Motion Predictor. Soft mAP is the official ranking metric, while miss rate is the secondary ranking metric. The first place is denoted by bold, the second place by underline, and the third place by *asterisk.
@@ IV-C Robustness
Unlike other models(see Fig.7), which exhibit a significant degradation in accuracy as the missing observation ratios increase, our approach proves to be more robust and reliable, particularly in challenging scenarios with higher levels of data loss. The inclusion of the recovery module demonstrates a clear improvement in performance compared to models like SIMPL and HIVT that lack this feature. When the recovery method is applied, these models consistently achieve lower brier-minFDE and minADE values, even as the percentage of missing observations increases (see Fig. 8). We also performed the same study using MTR as the baseline model on the Waymo motion dataset, evaluating the SoftmAP and minADE metrics (see Fig. 9). This indicates that our approach effectively handles data loss, resulting in more reliable and stable predictions even in highly uncertain environments. Please check out our submission video (due to page limitation) for qualitative results demonstrating the robustness of our method against noisy input.

### v2

@@ Abstract
We introduce RMP-YOLO, a unified framework designed to provide robust motion predictions even with incomplete input data. Our key insight stems from the observation that complete and reliable historical trajectory data plays a pivotal role in ensuring accurate motion prediction. Therefore, we propose a new paradigm that prioritizes the reconstruction of intact historical trajectories before feeding them into the prediction modules. Our approach introduces a novel scene tokenization module to enhance the extraction and fusion of spatial and temporal features. Following this, our proposed recovery module reconstructs agents’ incomplete historical trajectories by leveraging local map topology and interactions with nearby agents. The reconstructed, clean historical data is then integrated into the downstream prediction modules. Our framework is able to effectively handle missing data of varying lengths and remains robust against observation noise while maintaining high prediction accuracy. Furthermore, our recovery module is compatible with existing prediction models, ensuring seamless integration. Extensive experiments validate the effectiveness of our approach, and deployment in real-world autonomous vehicles confirms its practical utility. In the 2024 Waymo Motion Prediction Competition, our method, RMP-YOLO, achieves state-of-the-art performance, securing third place. Our code is open-source at https://github.com/ggosjw/RMP-YOLO.
@@ I INTRODUCTION
    Our method won third place in the 2024 Waymo Motion Prediction Competition. We deploy our algorithms on real vehicles to validate the effectiveness of our methods.
@@ II-A Motion Prediction
Recent advancements[7, 8, 9, 10, 11, 12] in multi-agent motion prediction have introduced several innovative approaches aimed at improving prediction accuracy and efficiency. Among these, two of the most influential prediction paradigms are MTR [7] and QCNet [8]. MTR addresses the trade-off between goal-based prediction[13, 14, 15, 16] and direct regression-based[17, 11, 15, 18, 19] prediction, by its proposed motion query pairs, while QCNet proposes a symmetric[20], translation-invarnt[21] input representation which enables reusable stream prediction. Plenty of derivative works have expanded on these foundational approaches, including EDA [22], MTR++ [9], ControlMTR [23], SIMPL [17], MGTR [24], and others [25, 26, 27, 28, 29]. However, a common limitation shared by these methods is that they are typically trained on nearly complete input trajectories. As a result, their performance may degrade when confronted with imperfect or incomplete data during inference (see Fig.7). To address this, we propose the RMP-YOLO framework, building upon MTR, to effectively handle input data imperfections while maintaining high prediction accuracy.
@@ II-B Imperfect Data Recovery
For feature distillation method[33], a well-trained teacher model using complete input trajectories guides a student mo
@@
del using incomplete input trajectories. POP[30] combines both methods to address partially observable prediction. While effective, this approach is complex, computationally expensive, and relies heavily on the quality of the teacher model. Some studies [34, 35] have investigated enhancing the robustness of predictors against input noise from the perspective of data poisoning. However, these approaches often overlook the impact of incomplete trajectories. Unlike previous methods, we propose a unified framework that explicitly recovers clean historical trajectories in the encoder stage and integrates this new information back into the agent tokens for further information fusion. As a result, our approach is robust to both noisy and incomplete data while maintaining excellent prediction performance. Additionally, our proposed recovery module is lightweight and plug-and-play.
@@ III-A Problem Formulation and Input Representation
Following MTR, we adopt a vectorized representation for both maps and agents. And our focus is on marginal motion prediction. The historical trajectories of ${N_{a}}$ traffic participants are denoted as $\mathcal{A}=\{a_{1},a_{2},\dots,a_{N_{a}}\}$, where each agent $a_{i}\in\mathbb{R}^{T_{p}\times F_{1}}$ has $T_{p}$ past timesteps and $F_{1}$ feature dimensions. The corresponding map is partitioned into $N_{l}$ polylines =$\mathcal{M}=\{m_{1},m_{2},\dots,m_{N_{l}}\}$, with each polyline $m_{i}\in\mathbb{R}^{N_{p}\times F_{2}}$ comprising $N_{p}$ points and $F_{2}$ feature dimensions. The predictor will anticipate $K$ different modality future trajectories $\mathcal{Z}=\{z_{1},z_{2},\dots,z_{N_{a}}\}$ over the future $T_{f}$ timesteps, where $z_{i}=\{z_{i}^{1},z_{i}^{2},\dots,z_{i}^{K}\}\in\mathbb{R}^{K\times T_{f}\times D}$. The confidence score for $z_{i}$ are denoted as $p_{i}=\{p_{i}^{1},p_{i}^{2},\dots,p_{i}^{K}\}$. Then for selected agent $z_{i}$, existing motion prediction task aims to estimate the distribution:
@@ III-B Network Encoder
where $\text{MHA}(\cdot,\cdot,\cdot)$ represents the multi-head attention function, $Q^{0}=[\mathcal{M}_{map},\mathcal{A}_{agent}]\in\mathbb{R}^{(N_{a}+N_{m})\times D_{1}}$, and $\mathcal{K}(\cdot)$ denotes the $K$-nearest neighbours (KNN) algorithm, which is used to identify the $K$ nearest tokens relative to each query. The term $\text{PE}(\cdot)$ refers to the sinusoidal positional encoding assigned to input tokens, incorporating the most recent position of each agent and the central point of each map polyline. The final output of the local attention layer will be sent into the MTR decoder, $[\mathcal{M}_{map},\mathcal{A}_{agent}]=Q^{Final}$.
@@ IV Experiments
TABLE I: Prediction on the test leaderboard of the motion prediction track of the Waymo Open Dataset Challenge. Our approach is termed RMP, i.e., Robust Motion Predictor. Soft mAP is the official ranking metric, while miss rate is the secondary ranking metric. The first place is denoted by bold, the second place by underline, and the third place by *asterisk.
@@ IV-C Robustness
Unlike other models(see Fig.7), which exhibit a significant degradation in accuracy as the missing observation ratios increase, our approach proves to be more robust and reliable, particularly in challenging scenarios with higher levels of data loss. The inclusion of the recovery module demonstrates a clear improvement in performance compared to models like SIMPL and HIVT which lack this feature. When the recovery method is applied, these models consistently achieve lower brier-minFDE and minADE values, even as the percentage of missing observations increases (see Fig. 8). We also performed the same study using MTR as the baseline model on the Waymo motion dataset, evaluating the SoftmAP and minADE metrics (see Fig. 9). This indicates that our approach effectively handles data loss, resulting in more reliable and stable predictions even in highly uncertain environments. Please check out our submission video (due to page limitation) for qualitative results demonstrating the robustness of our method against noisy input.